### v2

@@ 7 Conclusion
An in-depth examination of ion dynamics associated with ultrafast laser ablation plumes, from $5.2\times 10^{15}$ $W/cm^{2}$ pulses, emanating from nickel and a nickel/copper alloy surface provide detailed information about self-generated magnetic centrifuge fields in these ultrafast ablation plasmas. We examined, in particular, the effects of those fields on the radial mass enrichments around an axis normal to the ablating surface. It is shown that these energetic ablative ions, being subjected to a cylindrically symmetric longitudinal magnetic field which is oriented normal to the ablating surface, respond to a centrifuge mechanism that drives the radial distributions. Magnetic fields with peak values of 53 MG and average fields of 20 MG are implied from isotope enrichment data with local azimuthal electric currents on the order of $1.3\times 10^{6}$ amperes driving these fields.
@@
The analysis of isotope enrichment and our subsequent findings proceeded as follows: (1). Obtain mass separation ratios by casting the standard centrifuge equation into the initial geometry of an ablating cylindrical model that, after a short distance, transitions to an expanding cone. Apply this model of an expanding cone to find angular distributions of observed mass enrichments in terms of separation ratios. (2). These separation ratios are fitted to the centrifuge equation providing an overall averaged angular rotation rate for ions of $3.2\times 10^{9}$ radians/second. (3). The angular data scans are mapped back to the initial cylinder radial coordinate that allows detailed examination of the cyclotron orbits of ions at specific cylinder radii resulting in specific ion rotation rates at those radii. These are then converted to local magnetic field values, through the cyclotron equation, that provides a radial magnetic field distribution within the cylindrical centrifuge. This field distribution then provides a mathematical model that is used to extract current densities and enveloping currents responsible for the magnetic centrifuge mechanism. (4). The magnetic fields and currents are then used to construct the rigid rotor model of plasma rotation resulting in a value of $2.59\times 10^{5}$ radians per second, significantly lower than the spatially averaged cyclotron rates. (5). Anomalous localized charge state enrichments are examined in terms of resonant energy absorption from plasma waves that are identified as Ion Bernstein Waves (IBW). These are found to provide an excellent fit to the observed resonance data. A time dependent radial magnetic field profile is found necessary to properly fit these resonant data points. It is determined that this initial magnetic field lasts for at least 885 picoseconds and thereafter evolves to a stable quasi equilibrium configuration that lasts for as long as 4.85 nanoseconds before dissipating.
@@ 8.1 Coulomb Explosion
As explained earlier, the ablation process under femtosecond laser ablation occurs by a coulomb explosion due to the lighter electrons being stripped from the surface leaving a high concentration of positive ions in a solid density plasma condition. We described this explosion, in a previous report in terms of a parallel capacitor electric field model VanRompay et al. (1998). That model predicts a linear relationship between the charge state and its energy. The data below show that a quadratic behavior is more typical and needs to be analyzed in those terms. See Figure 4 for details where the data points in Figure 9 are taken as the most probable energy at each charge state.
@@ 8.2 Collimation Restrictions
The sector analyzer has a 2 mm acceptance aperture at its input Van Rompay (2003); VanRompay et al. (2000). The angular resolution $\Delta\theta$ of this ion detector is the half-angle of the acceptance cone subtended by the detector aperture as seen from the target:
@@ 8.3 Conductivity Perpendicular to B
Begin with the steady-state force balance equation for electrons in a magnetized, collisional plasma Boyd and Sanderson (2003); Bittencourt (2013):
@@
This result shows that the presence of a magnetic field suppresses the ability of electrons to carry current across field lines. In the limit of strong magnetization ($\omega_{ce}\gg\nu_{ei}$), the perpendicular conductivity becomes significantly smaller than the un-magnetized case. In contrast, the conductivity parallel to the field remains unaffected and equals $\sigma_{0}$. It is seen that the components of the tensor are controlled by the electron cyclotron frequency and the electron-ion collision rate. These are primarily controlled by the electron density, electron plasma temperature, and magnetic field strength.
@@
Using the empirical formula for the electron-ion collision frequency Boyd and Sanderson (2003) (with $T_{e}$ in eV, $n_{e}$ in m-3):
@@ Acknowledgments
The authors would like to thank the University of Michigan and Gerard Mourou, while director at the Center for Ultrafast Optical Science, for providing support and resources during the experimental phase of this work. The equipment used to collect these data was assembled through a grant from the Air Force Office of Scientific Research.

### v4

@@ 8 Conclusion
An in-depth examination of ion dynamics associated with ultrafast laser ablation plumes, from $5.2\times 10^{15}$ $W/cm^{2}$ pulses, emanating from nickel and a nickel/copper alloy surface provide detailed information about self-generated magnetic centrifuge fields in these ultrafast ablation plasmas. We examined, in particular, the effects of those fields on the radial mass enrichments around an axis normal to the ablating surface. It is shown that these energetic ablative ions, being subjected to a cylindrically symmetric longitudinal magnetic field which is oriented normal to the ablating surface, respond to a centrifuge mechanism that drives the radial distributions. Magnetic fields with peak effective value of 53 MG and average effective fields of 20 MG are implied from isotope enrichment data with effective local azimuthal electric currents on the order of $1.3\times 10^{6}$ amperes driving the fields.
These effective field values are further analyzed in terms of contributions from a longitudinal $B_{z}$ field plus an equivalent contribution from IBW electrostatic waves that greatly enhance the observed isotope enrichment and inflate these derived field values.
@@
The analysis of our data, and our subsequent findings, proceeded as follows: (1). Obtain mass separation ratios by casting the standard centrifuge equation into the initial geometry of an ablating cylindrical model that, after a short distance, transitions to an expanding cone. Apply this model of an expanding cone to find angular distributions of observed mass enrichments in terms of separation ratios. (2).These separation ratios are fitted to the centrifuge equation providing an overall averaged angular rotation rate for ions with an effective value of $3.2\times 10^{9}$ radians/second. (3).The angular data scans are mapped back to the initial cylinder radial coordinate that allows detailed examination of the cyclotron orbits of ions at specific cylinder radii resulting in ion rotation rates at those radii. These are then converted to local magnetic field values, through the cyclotron equation, that provides a radial magnetic field distribution within the cylindrical centrifuge. This field distribution then provides a mathematical model that is used to extract current densities and enveloping currents responsible for the magnetic centrifuge mechanism. (4). The magnetic fields and currents are then used to construct the rigid rotor model of plasma rotation resulting in a value of $2.59\times 10^{5}$ radians per second, significantly lower than the spatially averaged cyclotron rates. (5). Anomalous localized charge state enrichments are examined in terms of resonant energy absorption from plasma waves that are identified as Ion Bernstein Waves (IBW). These are found to provide an excellent fit to the observed resonance data. A time dependent radial magnetic field profile is found necessary to properly fit these resonant data points. It is determined that this initial magnetic field lasts for at least 885 picoseconds and thereafter evolves to a stable quasi equilibrium configuration that lasts for as long as 4.85 nanoseconds before dissipating.
[…]
This work has relevance to applications in isotope harvesting and provides insight into laser plasma physics relative to direct ignition fusion research, especially the proton-Boron 11 programs. In addition it demonstrates how isotopic analysis of laser plasmas can be used to derive a fuller understating of ablation phenomena, in particular its use in isotopically enriched thin film and nano-cluster deposition.
@@ 9.1 Coulomb Explosion
As explained earlier, the ablation process under femtosecond laser ablation occurs by a coulomb explosion due to the lighter electrons being stripped from the surface leaving a high concentration of positive ions in a solid density plasma condition. We described this explosion, in a previous report in terms of a parallel capacitor electric field model vanrompay1998pulse. That model predicts a linear relationship between the charge state and its energy. The data below show that a quadratic behavior is more typical and needs to be analyzed in those terms. See Figure 4 for details where the data points in Figure 9 are taken as the most probable energy at each charge state.
@@ 9.2 Collimation Restrictions
The sector analyzer has a 2 mm acceptance aperture at its input van2003mass; vanrompay2000isotope. The angular resolution $\Delta\theta$ of this ion detector is the half-angle of the acceptance cone subtended by the detector aperture as seen from the target:
@@ 9.3 Conductivity Perpendicular to B
Begin with the steady-state force balance equation for electrons in a magnetized, collisional plasma boyd2003physics; bittencourt2013fundamentals:
@@
This result shows that the presence of a magnetic field suppresses the ability of electrons to carry current across field lines. In the limit of strong magnetization ($\omega_{ce}\gg\nu_{ei}$), the perpendicular conductivity becomes significantly smaller than the un-magnetized case. In contrast, the conductivity parallel to the field remains unaffected and equals $\sigma_{0}$. It is seen that the components of the tensor are controlled by the electron cyclotron frequency and the electron-ion collision rate. These are, in turn, controlled by the electron density, electron plasma temperature, and magnetic field strength.
@@
Using the empirical formula for the electron-ion collision frequency boyd2003physics (with $T_{e}$ in eV, $n_{e}$ in m-3):
@@ Acknowledgments
The authors would like to thank the University of Michigan and Gerard Mourou, while he was director at the Center for Ultrafast Optical Science, for providing support and resources during the experimental phase of this work. The equipment used to collect these data was assembled through a grant from the Air Force Office of Scientific Research.